OpenStudy (anonymous):

Lenny runs a 100-metercourse in 35 seconds.Gary runs a 450-meter course in 112.5 seconds.Bruford runs 950-meter course in 237.5 seconds.Which athelete is the fastest?Round each speed to one decimal place. A. Lenny B. Gary C. Bruford D. They each travel at the same average speed. Which one?PLEASE!!!!!!!!!!!!!!

OpenStudy (anonymous):

simple , find out the speed of all of them per second , Lenny = 100/35 Gary = 450/112.5 Bruford = 950/237.5

OpenStudy (anonymous):

Tell me if u have any probs

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Lenny = 2.85 = 2.8m/s Gary = 4m/s Bruford = 4m/s